### v1

# Audio dequantization using instantaneous frequency

Abstract
We present a dequantization method that employs a phase-aware regularizer, originally successfully applied in an audio inpainting problem. The method maintains the temporal continuity of sinusoidal components in the audio signal time-frequency representation and avoids the energy loss artifacts commonly encountered with $\ell_{1}$-based regularization approaches. The proposed method is called the Phase-Aware Audio Dequantizer (PHADQ). The methods are evaluated using the objective metric SDR and PEMO-Q ODG.
Index Terms—  Instantaneous frequency, audio, dequantization, bit depth, sparsity.
1 Introduction
Quantization is a nonlinear distortion that arises from rounding during the A/D conversion or bit-depth reduction. This distortion appears as the quantization noise, which, at low bit depths (less than 16 bits per sample), results in unpleasant auditory artifacts [1]. Dequantization is the attempt to recover a signal close to its original form, based on the observed quantized samples. The problem is ill-conditioned; therefore, reconstruction involves a form of regularization. For instance, is the application of sparsity [2] is one potential approach. Autoregressive modeling is another modeling approach to dequantization [3, 4, 5].
In this paper, we present a dequantization method that employs a phase-aware regularizer introduced in [6], originally successfully applied in an audio inpainting problem. The method maintains the temporal continuity of sinusoidal components in the audio signal time–frequency representation and avoids the energy loss artifacts commonly encountered with $\ell_{1}$-based regularization approaches [6, 7].
2 Review of PHAIN (Phase-aware audio inpainter)
Audio inpainting is the task of recovering portions of an audio signal that are considered lost or unreliable [8, 9, 10, 7]. Let ${\mathbf{y}}\in\mathbb{R}^{L}$ be an audio signal with unreliable parts. PHAIN [6] aims to reconstruct $\mathbf{y}$} using the instantaneous frequency of the signal by solving the following optimization problem:
$\hat{{\mathbf{x}}}=\operatorname*{arg\,min}\limits_{{\mathbf{x}}\in\mathbb{R}^{L}}\ \ \lambda\|DR_{\omega_{\mathbf{s}}}G_{g}{\mathbf{x}}\|_{1}+\imath_{\mathit{\Gamma}}({\mathbf{x}}).$ (1)
Problem (1) contains a penalty $\|DR_{\omega_{\mathbf{s}}}G_{g}{\mathbf{x}}\|_{1}$ weighted by a positive $\lambda$, and an indicator function $\imath_{\mathit{\Gamma}}({\mathbf{x}})$ taking value $\infty$ when ${\mathbf{x}}\notin\Gamma$, thus enforcing the estimation to retain the reliable parts of the signal corresponding to the set of feasible signals $\Gamma$. The linear operators in the penalty function are:
  • The discrete Gabor transform (DGT) $G$ with a window $g$.
  • The phase correction $R$ of a spectrogram, based on the time derivative of the phase $\omega_{\mathbf{s}}$ (i.e., the instantaneous frequency) of a signal $\mathbf{s}$. Phase correction is performed using the cumulative sum of $\omega_{\mathbf{s}}$ as follows:
    $(R_{\omega_{\mathbf{s}}}\mathbf{z})[m,n]=\text{e}^{-2\pi\text{i}a\sum_{t=0}^{n-1}\omega_{\mathbf{s}}[m,t]/M}z[m,n],$ (2)
    where $a$ denotes the hop-size of the Gabor transform, and $n,m$ represent the time and frequency indices, respectively.
  • The time-directional first-order difference
    $(D\mathbf{z})[m,n]=z[m,n]-z[m,n+1].$ (3)
The convex $\ell_{1}$-norm denoted $\|\cdot\|_{1}$ quantifies time-directional variation of the phase-corrected time-frequency representation of $\mathbf{x}$}. Therefore, when used as a penalty, it promotes sinusoidal components and, conversely, attenuating components whose phase cannot be adequately approximated as a linear function, such as noise or transients. As a consequence, PHAIN to a great extent solves the problem encountered in plain sparsity-based methods which is loosing signal energy inside the gap [7]. When $\omega_{\mathbf{s}}$ is fixed, the problem (1) is convex and solvable via the Chambolle–Pock algorithm [11, 12].
Based on the regime of treating the instantaneous frequency during the reconstruction, PHAIN can further be classified as follows:
  • B-PHAIN, which uses the instantaneous frequency $\omega_{\mathbf{y}}$ of the degraded signal including its unreliable parts.
  • U-PHAIN, which updates the instantaneous frequency according to the current estimate of $\hat{{\mathbf{x}}}$.
3 Method
The call the proposed method simply the Phase-Aware Audio Dequantizer (PHADQ). In the method, we employ the same penalty as in PHAIN, but for a different restoration task.
Let ${\mathbf{y}}^{\text{q}}\in\mathbb{R}^{L}$ be a quantized observation of an original audio signal with a quantization step $\Delta$. When employing uniform quantization, the quantization step size is given by $\Delta=2^{-w+1}$, where $w$ denotes the word length in bits per sample (bps). We recover the signal by solving the same convex problem as in (1), except that no reliable parts of the signal are available, and therefore the set of feasible signals is now characterized by:
$\Gamma=\{{\mathbf{y}}\in\mathbb{R}^{L}\ |\ \|{\mathbf{y}}-{\mathbf{y}}^{\text{q}}\|_{\infty}<\Delta/2\},$ (4)
which means that a feasible sample must lie within the quantization interval corresponding to the original signal.
Using the indicator function $\imath_{\mathit{\Gamma}}$, we obtain the so-called consistent variant of the problem solution. An alternative approach is to relax the constraint defined by the indicator function and employ the (euclidean) distance to the set of feasible samples $d_{\Gamma}$. Problem (1) is then transformed to its inconsistent variant:
$\hat{{\mathbf{x}}}=\operatorname*{arg\,min}\limits_{{\mathbf{x}}\in\mathbb{R}^{L}}\ \ \lambda\|DR_{\omega_{\mathbf{s}}}G_{g}{\mathbf{x}}\|_{1}+\cfrac{1}{2}d^{2}_{\Gamma}({\mathbf{x}}).$ (5)
By choosing to update or not to update the instantaneous frequency estimate, we distinguish between the B-PHADQ and U-PHADQ, respectively.
3.1 B-PHADQ
The consistent and inconsistent problems (problems (1) and (5), respectively) are solved by Algorithm 1, which constitutes a specific instance of the Chambolle–Pock algorithm, wherein ${\mathbf{x}}\in\mathbb{R}^{L}$, ${\mathbf{p}}\in\mathbb{R}^{L}$ is a primal variable, ${\mathbf{q}}\in\mathbb{C}^{P}$ is a dual variable and ${\mathbf{u}}\in\mathbb{R}^{L}$ is a temporary variable. Positive scalars $\tau$, $\sigma$ are the step sizes and $\rho$ is the relaxation factor [12]. The scalar $\lambda>0$ is the weighting factor for the penalty function, having also an effect on the reconstruction quality and the convergence speed. Before applying the algorithm, we need to compute the instantaneous frequency $\omega_{\mathbf{s}}$, which in the case of B-PHADQ corresponds to the quantized observation, following [6]:
$\omega_{\mathbf{s}}[m,n]=\omega_{\mathbf{y}^{\text{q}}}[m,n]=-\text{Im}\Bigg[\cfrac{(G_{g^{\prime}}{\mathbf{y}^{\text{q}}})[m,n]}{(G_{g}{\mathbf{y}^{\text{q}}})[m,n]}\Bigg],$ (6)
where $G$ denotes the Gabor transform, $g^{\prime}$ is the time derivative of the window function, and $\text{Im}[\cdot]$ denotes the imaginary part.
Choose parameters $\tau,\sigma,>0,\ \rho\in[0,1]$ and initial values ${\mathbf{x}}^{(0)},{\mathbf{p}}^{(0)},{\mathbf{q}}^{(0)}$, initialize $\omega_{\mathbf{s}}$ by (6)
for $i=0,1,\dots$ do
 ${\mathbf{q}}^{(i+1)}=\text{clip}_{\lambda}({\mathbf{q}}^{(i)}+\sigma DR_{\omega_{\mathbf{s}}}G_{g}{\mathbf{x}}^{(i)})$
 ${\mathbf{u}}={\mathbf{p}}^{(i)}-\tau G^{*}_{g}R^{*}_{\omega_{\mathbf{s}}}D^{*}{\mathbf{q}}^{(i+1)}$   % auxiliary
 ${\mathbf{p}}^{(i+1)}=\text{proj}_{\Gamma}({\mathbf{u}})$   % consistent
 ${\mathbf{p}}^{(i+1)}=\frac{1}{\tau+1}(\tau\ \text{proj}_{\Gamma}({\mathbf{u}})+{\mathbf{u}})$   % inconsistent
 ${\mathbf{x}}^{(i+1)}={\mathbf{p}}^{(i+1)}+\rho({\mathbf{p}}^{(i+1)}-{\mathbf{p}}^{(i)})$
end for
return ${\mathbf{p}}^{(i+1)}$
Algorithm 1 B-PHADQ
The algorithm utilizes proximal operators of the functions present in the minimization, which in the case of the $\ell_{1}$ norm correspond to the soft-thresholding, in the case of the consistent indicator function it is the projection onto the set $\Gamma$. For the inconsistent case it is given by
$\text{prox}_{\alpha d^{2}_{\Gamma}/2}({\mathbf{x}})=\cfrac{1}{1+\alpha}\big(\alpha\text{proj}_{\Gamma}({\mathbf{x}})+{\mathbf{x}}\big).$ (7)
The asterisk denotes the adjoint of a linear operator. In the algorithm 1, the term ${\mathbf{p}}^{(i+1)}$ is defined differently for the consistent and inconsistent variant. Finally, $\text{clip}_{\lambda}$ is defined as $\text{clip}_{\lambda}=\text{Id}-\text{soft}_{\lambda}$.
3.2 U-PHADQ
As in U-PHAIN [6], an alternative approach to problems (1) and (5) is to regularly update the instantaneous frequency employed in the algorithm 1. This is done after a prescribed number of iterations. The new instantaneous frequency is calculated for an estimate $\mathbf{p}$}, rather than $\mathbf{y}^{\text{q}}$, according to the current reconstruction. U-PHADQ therefore contains Algorithm 1 within an additional loop, in the same way as in U-PHAIN [6].
4 Experiment and Results
For the experiment, we selected musical recordings from the IRMAS [^1] https://www.upf.edu/web/mtg/irmas and EBU SQAM [^2] https://tech.ebu.ch/publications/sqamcd datasets, and these selections are evaluated separately. The same datasets were used to evaluate the performance of the PHAIN method for audio inpainting [6].
From the IRMAS database, we selected 50 excerpts from various genres, featuring mixtures of different musical instruments (guitar, piano, violin, and vocals) included in the IRMAS-TestingData. The recordings have a sampling frequency of 44.1 kHz. Since they are originally in stereo, only the first channel was used, and each excerpt was truncated to 7 seconds. From the EBU SQAM database, we selected 10 recordings, mostly of solo instruments. The recordings have a sampling frequency of 44.1 kHz, and segments of approximately 6 seconds were used.
Audio has been peak-normalized before any processing to make the most of the available dynamic range. The audio was quantized to seven different word lengths, ranging from 2 to 8 bps using the mid-riser uniform quantization. In the experiment, the discrete Gabor transform was used with a Hann window of 8,192 samples in length, 75% overlap and 16,384 frequency channels. All computation was performed in MATLAB R2025a.
The parameters of Algorithm 1 were manually tuned. We used $\tau=1,\ \sigma=1$ and $\rho=1/3$. For better results, the balancing parameter $\lambda$ of the task needs to be set depending on the word length; we set it as shown in Table 1.
Table 1: Set of $\lambda$ according to word length
bps 2 3 4 5 6 7 8
$\lambda$ $10^{-1}$ $10^{-1}$ $10^{-3}$ $10^{-2}$ $10^{-3}$ $10^{-4}$ $10^{-4}$
In the experiments, the performance of the methods was evaluated using the objective metric SDR, which expresses the signal-to-distortion ratio and measures the physical similarity of waveforms. We also employed the PEMO-Q objective difference grade (ODG) [13], which considers the perceptual characteristics of the audio signals. The ODG is measured on a scale from 4-4 to 0 (worst to best).
As a baseline for comparison, we chose a sparsity-based dequantization method, namely the Chambolle–Pock algorithm (CP) performing best in the study [2]. This method solves the optimization problem
$\hat{{\mathbf{x}}}=\operatorname*{arg\,min}\limits_{{\mathbf{x}}\in\mathbb{R}^{L}}\ \ \lambda\|G_{g}{\mathbf{x}}\|_{1}+\imath_{\mathit{\Gamma}}({\mathbf{x}}),$ (8)
seeking for an analysis-sparse time-domain signal which fits the dequantization constraints. We used the Matlab code from [2] with the same DGT settings as in the PHADQ experiment.
When comparing the U-PHADQ and B-PHADQ variants, it was found that updating the instantaneous frequency within the course of iterations of U-PHADQ does not lead to better results in terms of the SDR. For this reason, in the experiments we focus only on the B-PHADQ variant. In addition, to see the maximum achievable performance of PHADQ, there is the B-PHADQ (oracle) variant, based on the instantaneous frequency of the original signal, which cannot be accessed in practical scenarios. For the oracle variant the consistent B-PHADQ was used. An example showing the evolution of SDR across iterations of PHADQ variants is presented in Fig. 1.
The results of the PHADQ method for both datasets show the best achievable SDR and ODG from a maximum of 200 iterations. The CP results were obtained after 500 iterations of the algorithm, achieving optimal ODG values [2].
Fig. 1: Reconstruction quality of a violin recording (6 bps word length) across iterations in terms of SDR. For U-PHADQ, $\omega_{\mathbf{s}}$ is updated every 10 iterations.
4.1 IRMAS dataset results
The SDR and ODG results for the IRMAS dataset are shown in Fig. 2. The reference B-PHADQ (oracle) shows the best results in both SDR and ODG, as expected. CP outperforms the proposed methods in terms of SDR. However, the consistent B-PHADQ performs no worse than by 1.3 dB across all word lengths. In terms of ODG, PHADQ proved to be superior, especially at a word length of 8 bps, where both consistent and inconsistent B-PHADQ exceeded CP by more than 0.3. Fig. 3 depicts the iteration number resulting in the best SDR and ODG reconstruction. At a word length of 8 bps, both B-PHADQ variants required no more than 60 algorithm iterations on average.
Fig. 2: Average SDR and ODG results on the IRMAS dataset
Fig. 3: Average number of iterations achieving the best SDR and ODG on the IRMAS dataset.
4.2 EBU SQAM dataset results
The results of the EBU SQAM dataset are presented Fig. 4 in the same way. There are some differences compared to the IRMAS dataset. The overall SDR is higher across all word lengths, whereas ODG shows lower values. This effect is due to the fact that the EBU SQAM dataset contains solo instruments, in which the (de)quantization error is hardly masked. In terms of SDR, CP again proves to be better than both consistent and inconsistent B-PHADQ. In terms of ODG, the B-PHADQ variants perform better than CP for most word lengths; however, the difference is not that pronounced than in the case of the IRMAS dataset.
Fig. 4: Average SDR and ODG on the EBU SQAM dataset
4.3 Computational considerations
For a 6-second-long excerpt, a single iteration of the consistent B-PHADQ variant takes about 0.12 second on a PC with AMD 3.80 GHz CPU and 32 GB RAM. Therefore, at a word length of 8 bps, optimal performance in terms of the ODG was reached in less than 7 seconds. A single iteration of the sparsity-based CP variant takes about 0.03 second, so a full reconstruction with 500 iterations requires approximately 15 seconds.
5 Conclusion
In this paper, we introduced a dequantization method that employs instantaneous frequency as a regularizer. The proposed approach demonstrated strong performance in objective evaluations, surpassing the reference method in terms of the perceptually-motivated ODG scale. Moreover, the method requires fewer algorithmic iterations than the sparsity-based CP algorithm, leading to reduced computational time. These findings highlight the potential of the proposed framework for efficient and high-quality restoration of quantized signals.
Codes for Matlab are publicly available. [^3] https://github.com/vojtechkovanda/PHADQ