### v2

# Audio dequantization using instantaneous frequency

Abstract
We present a dequantization method that employs a phase-aware regularizer, originally successfully applied in an audio inpainting problem. The method promotes a temporal continuity of sinusoidal components in time-frequency representation of the audio signal, and avoids energy loss artifacts commonly encountered with $\ell_{1}$-based regularization approaches. The proposed method is called the Phase-Aware Audio Dequantizer (PHADQ). The method are evaluated against the state-of-the-art using the SDR and PEMO-Q ODG objective metrics, and a subjective MUSHRA-like test.
Index Terms—  Instantaneous frequency, audio, dequantization, bit depth, sparsity.
1 Introduction
Quantization is a nonlinear distortion that arises from rounding during the A/D conversion or bit-depth reduction. This distortion appears as the quantization noise, which, at low bit depths (less than 16 bits per audio sample), results in unpleasant auditory artifacts [13]. Dequantization is the attempt to recover a signal close to its original form, based on the observed quantized samples. The problem is ill-conditioned; therefore, the reconstruction involves a form of regularization. For instance, is the application of sparsity [14] is one common approach. Autoregressive modeling is another modeling approach to dequantization [10, 12, 11].
In this paper, we present a dequantization method that employs a phase-aware regularizer introduced in [9], originally successfully applied in the context of the audio inpainting problem. The regularizer maintains the temporal continuity of sinusoidal components in the audio signal time-frequency representation and avoids the energy loss artifacts commonly encountered with $\ell_{1}$-based regularization approaches [9, 6]. The goal of this paper is to investigate whether the phase-aware regularization is beneficial also in the dequantization setup. The reconstruction quality and computational demands are compared to a baseline audio dequantization method [14].
2 Review of PHAIN (Phase-aware audio inpainter)
Audio inpainting is the task of recovering portions of an audio signal that are considered lost or unreliable [1, 5, 7, 6]. Let ${\mathbf{y}}\in\mathbb{R}^{L}$ be an audio signal with unreliable parts. PHAIN [9] aims to reconstruct $\mathbf{y}$} using the instantaneous frequency of the signal by solving the following optimization problem:
$\hat{{\mathbf{x}}}=\operatorname*{arg\,min}\limits_{{\mathbf{x}}\in\mathbb{R}^{L}}\ \ \lambda\|DR_{\omega_{\mathbf{s}}}G_{g}{\mathbf{x}}\|_{1}+\imath_{\mathit{\Gamma}}({\mathbf{x}}).$ (1)
Problem (1) contains a penalty $\|DR_{\omega_{\mathbf{s}}}G_{g}{\mathbf{x}}\|_{1}$ weighted by a positive $\lambda$, and an indicator function $\imath_{\mathit{\Gamma}}({\mathbf{x}})$ taking value $\infty$ when ${\mathbf{x}}\notin\Gamma$, thus enforcing the estimation to retain the reliable parts of the signal corresponding to the set of feasible signals $\mathit{\Gamma}$}. The linear operators in the penalty function are:
  • The discrete Gabor transform (DGT) $G_{g}$ employing a window $g$.
  • The phase correction $R$ of a spectrogram, based on the time derivative of the phase $\omega_{\mathbf{s}}$ (i.e., the instantaneous frequency) of a signal $\mathbf{s}$. Phase correction is performed using the cumulative sum of $\omega_{\mathbf{s}}$ as follows:
    $(R_{\omega_{\mathbf{s}}}\mathbf{z})[m,n]=\text{e}^{-2\pi\text{i}a\sum_{t=0}^{n-1}\omega_{\mathbf{s}}[m,t]/M}z[m,n],$ (2)
    where $a$ denotes the hop-size of the Gabor transform, and $n,m$ represent the time and frequency indices, respectively. The size of $\omega_{\mathbf{s}}$ is identical to the size of the spectrogram obtained by $G_{g}$.
  • The time-directional first-order difference
    $(D\mathbf{z})[m,n]=z[m,n]-z[m,n+1].$ (3)
The convex $\ell_{1}$-norm denoted $\|\cdot\|_{1}$ quantifies time-directional variation of the phase-corrected time-frequency representation of $\mathbf{x}$}. When used as a penalty, it promotes sinusoidal components and, conversely, attenuates components whose phase cannot be adequately approximated as a linear function, such as noise or transients. As a consequence, PHAIN to a great extent solves the problem encountered in plain sparsity-based methods which is loosing signal energy inside the gap [6]. When $\omega_{\mathbf{s}}$ is fixed, the problem (1) is convex and solvable via the Chambolle–Pock algorithm [2, 3].
Based on the regime of treating the instantaneous frequency during the process of reconstruction, PHAIN can further be classified as follows:
  • B-PHAIN, which uses the instantaneous frequency $\omega_{\mathbf{y}}$ of the degraded signal, i.e. including unreliable parts.
  • U-PHAIN, which updates the instantaneous frequency according to the estimate of $\hat{{\mathbf{x}}}$ during iterations.
3 Method
We call the proposed method simply the Phase-Aware Audio Dequantizer (PHADQ). In the method, we employ the same penalty as in PHAIN, but for a different restoration task.
Let ${\mathbf{y}}^{\text{q}}\in\mathbb{R}^{L}$ be a quantized observation of an original audio signal with a quantization step $\Delta$. When employing uniform quantization, the quantization step size is fixed and given by $\Delta=2^{-w+1}$, where $w$ denotes the word length in bits per sample (bps). We recover the signal by solving the same convex problem as in (1), except that now no reliable parts of the signal are available. Therefore, in PHADQ, the set of feasible signals is characterized by:
$\Gamma=\{{\mathbf{y}}\in\mathbb{R}^{L}\ |\ \|{\mathbf{y}}-{\mathbf{y}}^{\text{q}}\|_{\infty}<\Delta/2\},$ (4)
which means that a feasible sample must lie within the quantization interval corresponding to the original signal.
Using the indicator function $\imath_{\mathit{\Gamma}}$, we obtain the so-called consistent variant of the problem solution. An alternative approach is to relax the constraint defined by the indicator function and employ the (euclidean) distance to the set of feasible samples $d_{\Gamma}$. Problem (1) is then transformed to its inconsistent variant:
$\hat{{\mathbf{x}}}=\operatorname*{arg\,min}\limits_{{\mathbf{x}}\in\mathbb{R}^{L}}\ \ \lambda\|DR_{\omega_{\mathbf{s}}}G_{g}{\mathbf{x}}\|_{1}+\cfrac{1}{2}\,d^{2}_{\Gamma}({\mathbf{x}}).$ (5)
By choosing whether to update the instantaneous frequency estimate, we distinguish between the B-PHADQ and U-PHADQ, respectively.
3.1 B-PHADQ
Both the consistent and inconsistent restoration problems (i.e., (1) and (5), respectively) are solved by Algorithm 1, which constitutes a specific instance of the Chambolle–Pock algorithm, wherein ${\mathbf{x}}\in\mathbb{R}^{L}$, ${\mathbf{p}}\in\mathbb{R}^{L}$ is a primal variable, ${\mathbf{q}}\in\mathbb{C}^{P}$ is a dual variable and ${\mathbf{u}}\in\mathbb{R}^{L}$ is a temporary variable. Positive scalars $\tau$, $\sigma$ are the step sizes and $\rho$ is the relaxation factor [3]. The scalar $\lambda>0$ is the weighting factor for the penalty function, having also an effect on the reconstruction quality and the convergence speed. Before applying the algorithm, an initial estimate of the instantaneous frequency $\omega_{\mathbf{s}}$ has to be established, which corresponds to the following formula using the quantized observation [9]:
$\omega_{\mathbf{s}}[m,n]=\omega_{\mathbf{y}^{\text{q}}}[m,n]=-\text{Im}\Bigg[\cfrac{(G_{g^{\prime}}{\mathbf{y}^{\text{q}}})[m,n]}{(G_{g}{\mathbf{y}^{\text{q}}})[m,n]}\Bigg].$ (6)
Here, the Gabor transform is used twice, once with window function $g$ and the second time with $g^{\prime}$, the time derivative of $g$. $\text{Im}[\cdot]$ denotes the imaginary part.
Choose parameters $\tau,\sigma,>0,\ \rho\in[0,1]$ and initial values ${\mathbf{x}}^{(0)},{\mathbf{p}}^{(0)},{\mathbf{q}}^{(0)}$, initialize $\omega_{\mathbf{s}}$ by (6)
for $i=0,1,\dots$ do
 ${\mathbf{q}}^{(i+1)}=\text{clip}_{\lambda}({\mathbf{q}}^{(i)}+\sigma DR_{\omega_{\mathbf{s}}}G_{g}{\mathbf{x}}^{(i)})$
 ${\mathbf{u}}={\mathbf{p}}^{(i)}-\tau G^{*}_{g}R^{*}_{\omega_{\mathbf{s}}}D^{*}{\mathbf{q}}^{(i+1)}$   % auxiliary
 ${\mathbf{p}}^{(i+1)}=\text{proj}_{\Gamma}({\mathbf{u}})$   % consistent
 ${\mathbf{p}}^{(i+1)}=\frac{1}{\tau+1}(\tau\ \text{proj}_{\Gamma}({\mathbf{u}})+{\mathbf{u}})$   % inconsistent
 ${\mathbf{x}}^{(i+1)}={\mathbf{p}}^{(i+1)}+\rho({\mathbf{p}}^{(i+1)}-{\mathbf{p}}^{(i)})$
end for
return ${\mathbf{p}}^{(i+1)}$
Algorithm 1 B-PHADQ
The algorithm utilizes proximal operators of the functions present in the minimization, which in the case of the $\ell_{1}$ norm correspond to the soft-thresholding, and in the case of the consistent indicator function it is the projection onto the set $\Gamma$. For the inconsistent case it is given by
$\text{prox}_{\alpha d^{2}_{\Gamma}/2}({\mathbf{x}})=\cfrac{1}{1+\alpha}\big(\alpha\,\text{proj}_{\Gamma}({\mathbf{x}})+{\mathbf{x}}\big).$ (7)
The asterisk in Algorithm 1 denotes the adjoint of a linear operator. Note that the term ${\mathbf{p}}^{(i+1)}$ is computed differently for the consistent and for the inconsistent variant. Finally, $\text{clip}_{\lambda}$ is defined as $\text{clip}_{\lambda}({\mathbf{x}})={\mathbf{x}}-\text{soft}_{\lambda}({\mathbf{x}})$.
3.2 U-PHADQ
As in U-PHAIN [9], an alternative approach to problems (1) and (5) is to regularly update the instantaneous frequency that is fixed in B-PHADQ in Algorithm 1. In U-PHADQ, such an update is performed after a user-prescribed number of iterations of Algorithm 1. The new instantaneous frequency is calculated for $\mathbf{p}$} rather than from $\mathbf{y}^{\text{q}}$, this way corresponding to the current state of reconstruction. U-PHADQ therefore contains the main loop of Algorithm 1 as its inner loop, the same way as in U-PHAIN [9].
4 Experiment and Results
For the experiment, we selected musical recordings from the IRMAS [^1] https://www.upf.edu/web/mtg/irmas and EBU SQAM [^2] https://tech.ebu.ch/publications/sqamcd datasets, and these selections are evaluated separately. The same datasets were used to evaluate the performance of the PHAIN method for audio inpainting [9].
From the IRMAS database, we selected 50 excerpts from various genres, featuring mixtures of different musical instruments (guitar, piano, violin, and vocals) included in the IRMAS-TestingData. The recordings have a sampling frequency of 44.1 kHz. Since they are originally in stereo, only the first channel was used, and each excerpt was truncated to 7 seconds. From the EBU SQAM database, we selected 10 recordings, mostly of solo instruments. The recordings have a sampling frequency of 44.1 kHz, and segments of approximately 6 seconds were used.
Audio has been peak-normalized before any processing to make the most of the available dynamic range. The audio was quantized to seven different word lengths, ranging from 2 to 8 bps using the mid-riser uniform quantization. In the experiment, the Gabor transform with a Hann window of 8,192 samples in length, 75% overlap and 16,384 frequency channels has been used. All computation was performed in Matlab R2025a.
The parameters of Algorithm 1 were manually tuned. We used $\tau=1,\ \sigma=1$ and $\rho=1/3$. For better results, the balancing parameter $\lambda$ of the task needs to be set depending on the word length and variant; we set it as shown in Table 1. Indices c or i denote consistent or inconsistent variant. Note that $\lambda$ increases with decreasing bps, which is expected since the lower bps, the greater is the amount of error/noise. Also note that the inconsistent case requires lower $\lambda$ than its consistent counterpart, which makes sense, because in (5), the error is partially absorbed by the distance term, opposite to the hard constraint in (1).
Table 1: Values of $\lambda$ according to word length and method
bps 2 3 4 5 6 7 8
$\lambda_{\text{c}}$ $0.07$ $0.070$ $0.030$ $0.010$ $0.0010$ $0.0050$ $0.0002$
$\lambda_{\text{i}}$ $0.07$ $0.015$ $0.006$ $0.001$ $0.0008$ $0.0005$ $0.0002$
In the experiments, the performance of the methods was evaluated using the objective metric SDR, which expresses the signal-to-distortion ratio and measures the physical similarity of waveforms. We also employed the PEMO-Q objective difference grade (ODG) [4], which considers the perceptual characteristics of the audio signals. The ODG is measured on a scale from 4-4 to 0 (worst to best). To obtain subjective quality assessments, we employ the MUSHRA listening test [8].
As a baseline for comparison, we chose a sparsity-based dequantization method, namely the Chambolle–Pock algorithm (CP) performing best in the study [14]. This method solves the optimization problem
$\hat{{\mathbf{x}}}=\operatorname*{arg\,min}\limits_{{\mathbf{x}}\in\mathbb{R}^{L}}\ \ \lambda\|G_{g}{\mathbf{x}}\|_{1}+\imath_{\mathit{\Gamma}}({\mathbf{x}}),$ (8)
seeking for an analysis-sparse time-domain signal which fits the dequantization constraints. We used the Matlab code from [14] with the same DGT settings as in the PHADQ experiment.
We show an example showing the evolution of SDR across iterations of PHADQ variants presented in Fig. 1. When comparing the U-PHADQ and B-PHADQ variants, it was found that updating the instantaneous frequency within the course of iterations of U-PHADQ does not lead to better results in terms of the SDR. For this reason, in the experiments we focus only on the B-PHADQ variant. In addition, to see the maximum achievable performance of PHADQ, there is the B-PHADQ (oracle) variant, based on the instantaneous frequency of the original signal, which cannot be accessed in practical scenarios. For the oracle variant the consistent B-PHADQ was used.
The results of the PHADQ method for both datasets show ODG and SDR after 60 iterations of the algorithm, which represents a balance between the average number of iterations required to reach optimal SDR and ODG values. We note that a higher number of algorithm iterations does not guarantee improved ODG values; the suplementary graphs are publicly available [^3] https://vojtechkovanda.github.io/PHADQ/. The CP results were obtained after 500 iterations of the algorithm, achieving optimal ODG values [14].
Fig. 1: Reconstruction quality of a violin recording (6 bps word length) across iterations in terms of SDR. For U-PHADQ, $\omega_{\mathbf{s}}$ is updated every 10 iterations.
4.1 Objective evaluation
The SDR and ODG results for the EBU SQAM and IRMAS dataset are shown in Fig. 2 and Fig. 3 respectively. As expected, the reference B-PHADQ (oracle) exhibits the best results in ODG in both datasets. CP outperforms the proposed methods in terms of SDR. However, in terms of ODG, the consistent B-PHADQ outperforms CP across almost all word lengths in both datasets. In the EBU SQAM dataset, all methods achieve higher SDR but lower ODG. This is due to the fact that EBU SQAM contains recordings of solo instruments, making the differences between recordings more perceptible. On the other hand, the IRMAS dataset contains recordings with multiple instruments, which masks the quantization noise more effectively, resulting in higher ODG.
Fig. 2: Average SDR and ODG on the EBU SQAM dataset
Fig. 3: Average SDR and ODG results on the IRMAS dataset
4.2 Subjective evaluation
The subjective test was conducted on the EBU SQAM database, comparing the CP, consistent B-PHADQ, and inconsistent B-PHADQ methods. Only 6-bit and 7-bit word lengths were used in the test, as these are the highest bit depths at which quantization distortion is still relatively clearly perceptible. The test results are shown in Fig. 4 and confirm the objective evaluation.
Fig. 4: MUSHRA results for 6 and 7 bps quantization on EBU SQAM dataset
4.3 Computational considerations
For a 6-second-long excerpt, a single iteration of the consistent B-PHADQ variant takes about 0.07 second on a PC with AMD 3.80 GHz CPU and 32 GB RAM. Therefore a full, 60 iteration long reconstruction is obtained in 4.2 seconds. A single iteration of the sparsity-based CP variant takes about 0.03 second, thus a full reconstruction with 500 iterations requires approximately 15 seconds.
5 Conclusion
In this paper, we introduced a dequantization method that employs instantaneous frequency as a regularizer. The proposed approach demonstrated strong performance in objective evaluations. Moreover, the method requires fewer algorithmic iterations to converge than the sparsity-based CP algorithm, leading to reduced computational time. These findings highlight the potential of the proposed framework for efficient and high-quality restoration of quantized signals.
Codes for Matlab are publicly available. [^4] https://github.com/vojtechkovanda/PHADQ